A firm has total assets of $14 million and a debt/equity ratio of 0.75. Its sales are $10 million, and it has total fixed costs of $4 million. If the firm's EBIT is $2 million, its tax rate is 45%, and the interest rate on all of its debt is 10%, what is the firm's ROE?
Answer #1

The solution to the given question is provided below:
Debt=14*0.75/(1+0.75)
Equity=14/(1+0.75)
Net Income=(2-14*0.75/(1+0.75)*10%)*(1-45%)
RoE=(2-14*0.75/(1+0.75)*10%)*(1-45%)/(14/(1+0.75))=9.625%